### v1

# Real-Time Per-Garment Virtual Try-On with Temporal Consistency for Loose-Fitting Garments

Abstract.
Per-garment virtual try-on methods collect garment-specific datasets and train networks tailored to each garment to achieve superior results. However, these approaches often struggle with loose-fitting garments due to two key limitations: (1) They rely on human body semantic maps to align garments with the body, but these maps become unreliable when body contours are obscured by loose-fitting garments, resulting in degraded outcomes; (2) They train garment synthesis networks on a per-frame basis without utilizing temporal information, leading to noticeable jittering artifacts. To address these challenges, we propose a two-stage approach for robust semantic map estimation. First, we extract a garment-invariant representation from the raw input image. This representation is then passed through an auxiliary network to estimate the semantic map. This enhances the robustness of semantic map estimation under loose-fitting garments during garment-specific dataset generation. Furthermore, we introduce a recurrent garment synthesis framework that incorporates temporal dependencies to improve frame-to-frame coherence while maintaining real-time performance. We conducted qualitative and quantitative evaluations to demonstrate that our method outperforms existing approaches in both image quality and temporal coherence. Ablation studies further validate the effectiveness of the garment-invariant representation and the recurrent synthesis framework.
Virtual Try-On, Image Synthesis, Human-in-the-loop Machine Learning
copyright: noneconference: ; ; ccs: Computing methodologies Image processingccs: Human-centered computing Mixed / augmented realityccs: Computing methodologies Image-based rendering
((a)) Real-time virtual try-on
((b)) Loose-fitting garment try-on results
Figure 1. (a) Our virtual try-on system features an efficient inference pipeline that enables real-time performance. (b) Our method produces highly realistic $360^{\circ}$ virtual try-on results for loose-fitting garments with temporal consistency.
Figure 2. Illustration of the training and inference pipelines for applying the method proposed in (Wu et al., 2025) to virtually try on a loose-fitting garment. (a) During training, the garment synthesis network is trained using degraded DensePose (Güler et al., 2018) outputs. (b) During inference, a distribution mismatch between training and input data leads to suboptimal garment synthesis results.
1. Introduction
Virtual try-on technology enables customers to try on garments without physical access, has recently garnered significant attention from researchers in today’s e-commerce landscape. Existing image-based virtual try-on methods (Jetchev and Bergmann, 2017; Dong et al., 2019; Choi et al., 2021; Lee et al., 2022; Fang et al., 2024; Karras et al., 2024) require only a 2D in-shop image of the target garment to generate virtual try-on results, making them highly accessible. However, these approaches are predominantly tailored for retailers aiming to produce promotional images featuring fashion models, rather than for end-users seeking personalized virtual try-on experiences. This focus is clearly reflected in the biased datasets used for training, which are heavily skewed toward tall and slim fashion models. As a result, these methods fail to generalize well to the diverse body types of average consumers.
Unlike general image-based methods, per-garment virtual try-on methods (Chong et al., 2021; Wu et al., 2024, 2025) focus on customer-oriented virtual try-on by collecting garment-specific datasets and training garment-specific neural networks. Early per-garment methods (Chong et al., 2021; Wu et al., 2024) require an expensive robotic mannequin for dataset collection, making their results non-reproducible. (Wu et al., 2025) lower the barrier to dataset collection significantly by using real human bodies for dataset collection. This approach relies on DensePose (Güler et al., 2018) for estimating human body semantic maps, which are crucial for aligning the human body and the synthesized garment in 2D image space. However, it struggles with loose-fitting garments, as DensePose results degrade significantly for these types of clothing, causing a distribution discrepancy between the training and inference stages, as illustrated in Figure 2. Such discrepancies lead to degraded results during the inference stage, where users typically wear tight-fitting garments to facilitate accurate human body tracking.
Moreover, previous per-garment approaches rely solely on the current input frame for garment synthesis, neglecting the temporal continuity provided by preceding frames. As a result, their results often exhibit jittering artifacts—especially for loose-fitting garments, which naturally exhibit appearance variations even under identical poses.
To address these limitations, we introduce a garment-invariant representation that can be directly extracted from the raw input image. This representation is then transformed into a human body semantic map using an auxiliary network, enabling robust estimation of the human body semantic map under loose-fitting garments during garment-specific dataset generation. The garment-invariant representation leverages joint heatmaps and 3D human pose, both of which are inherently robust to garment variations. Furthermore, we propose a recurrent garment synthesis framework that ensures temporal consistency by conditioning the synthesis process on both the current input frame and the preceding temporal states.
Our recurrent garment synthesis framework is capable of processing input sequences of arbitrary length while maintaining high efficiency, achieving approximately 10 frames per second on a standard PC. This ensures a smooth and responsive virtual try-on experience for users, as demonstrated in Figure 1. Both qualitative and quantitative evaluations indicate that our method surpasses existing approaches in terms of visual quality and temporal consistency. We also conduct ablation studies to validate the necessity of the garment-invariant representation and the recurrent synthesis framework.
Our contributions can be summarized as follows:
  • We propose a garment-invariant representation for robust human body semantic map estimation under loose-fitting garments, enabling per-garment virtual try-on for such garments.
  • We develop a recurrent garment synthesis framework that supports real-time virtual try-on with temporal consistency, accommodating input sequences of arbitrary length.
  • We conduct comprehensive qualitative and quantitative comparisons with existing methods, along with ablation studies, to validate the effectiveness and necessity of each component in our proposed framework.
2. Related Work
2.1. 3D model-based Virtual Try-On
3D model-based virtual try-on methods (Guan et al., 2012; Sekine et al., 2014; Santesteban et al., 2019; Patel et al., 2020; Pan et al., 2022) typically represent garments using polygon meshes. These methods generate try-on results by simulating the interaction between clothing and a 3D human body model, leveraging either physics-based simulations (Cirio et al., 2014; Kaldor et al., 2008; Narain et al., 2012; Selle et al., 2008) or learning-based approaches (Casado-Elvira et al., 2022; Grigorev et al., 2023; Halimi et al., 2023; Lahner et al., 2018; Pan et al., 2022; Patel et al., 2020; Santesteban et al., 2019, 2022, 2021; Xiang et al., 2021). While these techniques can realistically capture garment deformations across diverse poses and viewpoints, they require labor-intensive manual 3D modeling for each garment, which poses a significant barrier to scalability. To address this, some recent approaches (Feng et al., 2022; Dong and Wang, 2023; Li et al., 2024; Lin et al., 2024; Chen et al., 2024a, b; Cao et al., 2024; He et al., 2025) adopt volumetric representations, such as Neural Radiance Fields (NeRF) (Mildenhall et al., 2021) and 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023), to model garments directly from monocular images. These representations facilitate automatic 3D garment reconstruction but are primarily designed for digital avatars and are not optimized for real-time virtual try-on experiences involving real human users.
In contrast, our method supports real-time virtual try-ons for real-world users and enables automatic pre-processing for each garment with minimal human intervention.
2.2. Image-based Virtual Try-On
Image-based virtual try-on methods eliminate the need for 3D modeling of garments, relying instead on a single 2D in-shop image to generate try-on results (Song et al., 2023). This makes them significantly more accessible and scalable compared to 3D model-based approaches. Several methods (Jetchev and Bergmann, 2017; Choi et al., 2021; Lee et al., 2022; Kim et al., 2024; Wang et al., 2024; Xu et al., 2024b) focus on producing static try-on images, which lack the ability to convey dynamic garment behavior. To address this, other approaches (Dong et al., 2019; Chen et al., 2021; Jiang et al., 2022; Fang et al., 2024; Xu et al., 2024a; Karras et al., 2024) incorporate inter-frame consistency to generate temporally coherent video try-on results. However, these methods typically suffer from high memory usage, restricting the length of video they can process. It is also important to note that most existing image-based virtual try-on systems are designed primarily for e-commerce retailers to showcase garments on idealized fashion models for marketing purposes. Their networks are typically trained on datasets biased toward tall and slim body types, resulting in suboptimal performance when applied to average users.
In contrast, our method is designed for customer-centric virtual try-ons and supports real-time virtual try-on with temporal consistency for arbitrary lengths of input frames.
2.3. Per-garment Virtual Try-On
Unlike image-based virtual try-on approaches that rely on universal networks trained across diverse garment types, per-garment methods focus on training garment-specific networks using dedicated datasets, resulting in superior try-on quality for specific garments. A pioneering effort by (Chong et al., 2021) introduced the use of a robotic mannequin to automate the collection of garment-specific datasets, enabling the training of garment-specific networks. However, this approach requires users to wear a physical measurement garment for accurate body tracking, which limits its practicality. To address this, (Wu et al., 2024) proposed a vision-based intermediate representation that eliminates the need for the physical measurement garment. More recently, (Wu et al., 2025) further reduced the barrier to dataset collection by leveraging real human bodies, thereby removing the dependency on costly robotic mannequins. Despite these advancements, (Wu et al., 2025) struggles with loose-fitting garments due to their reliance on human body semantic map estimation, which often fails under such garments, and the jittering artifacts resulting from a lack of inter-frame consistency enforcement.
Our method addresses these limitations by introducing a garment-invariant representation for human body semantic map estimation and a recurrent garment synthesis framework, enabling high-quality dataset generation and temporally consistent virtual try-ons for loose-fitting garments.
2.4. Domain Adaption
Domain adaptation aims to ensure the performance of machine learning models trained on a source domain when applied to a target domain with a different data distribution (Farahani et al., 2021). The predominant paradigm of domain adaptation is mapping the source and target domains into a shared space (Tzeng et al., 2019, 2017, 2015; Sankaranarayanan et al., 2018; Hoffman et al., 2016). (Hoffman et al., 2016) present a domain adaptation framework with fully convolutional networks for semantic segmentation. They learn the mapping to the shared space in an unsupervised manner by assuming source and target domains share the same label statistics. (Isola et al., 2017) directly learns the mapping from the source domain to the target domain in a supervised way. However, the correlations between the source and target domain are usually unknown. (Royer et al., 2020) and (Zhu et al., 2017) use cycle consistency to learn the mappings between unpaired source and target images. However, unpaired image-to-image translation struggles with high-resolution images, limiting their applicability in our scenario.
In this paper, we propose a shared representation space for images of humans wearing tight-fitting and loose-fitting garments. Our aim is to bridge the domain gap that causes existing human body semantic map estimation models, trained predominantly on tight-fitting garment images, to underperform when applied to loose-fitting garment images.
Figure 3. Overview of our method. We first train a per-person BodyMap network using images of a person wearing tight-fitting garments. This network is then used to robustly estimate body semantic maps for the same person in a loose-fitting garment. These estimated maps are subsequently used to train a per-garment ReGarSyn network. Once trained, the ReGarSynt network enables real-time virtual try-on for arbitrary persons.
Figure 4. DensePose (Güler et al., 2018) accurately estimates the human body semantic map when subjects wear tight-fitting garments (left). However, its performance significantly degrades with loose-fitting garments due to occlusions (right).
3. Method
3.1. Overview
Figure 3 provides an overview of our method. To mitigate the distribution gap between the training and inference stages, we aim to obtain robust human body semantic maps under loose-fitting garments. To achieve this, we first train a per-person BodyMap network using images of a person wearing arbitrary tight-fitting garments. This network can then be used to robustly estimate body semantic maps for the same person in the target loose-fitting garment by transforming the extracted garment-invariant representations into body semantic maps.
Subsequently, these estimated maps are used to train a per-garment Recurrent Garment Synthesis (ReGarSyn) network that learns the mapping from the input sequence to the target garment image sequence. Once trained, the ReGarSyn network enables real-time, temporally consistent virtual try-ons for arbitrary persons.
3.2. Design of Garment-Invariant Representation
(Wu et al., 2025) employ DensePose (Güler et al., 2018) for estimating the human body semantic map directly from the raw input image. However, DensePose varies significantly under loose-fitting garments, as shown in Figure 4. To ensure reliable human body semantic map estimation under loose-fitting garments, we need a garment-invariant representation that captures human body information and is invariant to the garment to serve as the intermediate representation. A natural candidate for such a representation is the set of human joints. However, not all joints are truly garment-invariant. The presence and type of garments can influence their detected positions and associated confidence scores, as shown in Figure 12. The human joint estimation results are generated by Sapiens (Khirodkar et al., 2024), which is the state-of-the-art human pose estimation method.
Although human joint estimation is not entirely garment-invariant, it can still serve as a robust representation by excluding joints sensitive to garment variations, such as the hips and knees. However, excluding these joints limits the expressiveness of the resulting pose representation. To address this limitation, we augment the remaining joint data with virtual measurement garment (Wu et al., 2025), which is a rendered image of the SMPL mesh that encapsulates the full 3D human pose. In contrast to (Wu et al., 2025), which uses an upper-body template mesh, we utilize a full-body template mesh to facilitate the representation of full-body pose.
As illustrated in Figure 5, given an input person image $\mathbf{I}\in\mathbb{R}^{3\times H\times W}$, we first estimate the 3D human pose using BEV (Sun et al., 2022), which outputs a SMPL mesh (Loper et al., 2015). To ensure garment invariance, we fix the body shape parameters that are sensitive to garment variations. Subsequently, we remove non-essential body parts (head, hands, and feet) and apply a grid-pattern texture to the remaining mesh. This textured mesh is then rendered into an RGB image $\mathbf{I}_{vm}\in\mathbb{R}^{3\times H\times W}$.
In parallel, we extract 133 joint heatmaps using Sapiens (Khirodkar et al., 2024). To further reduce sensitivity to garment variation, we discard the heatmaps corresponding to the hip and knee joints. The remaining heatmaps are grouped into three categories and compactly represented as an RGB image $\mathbf{I}_{shm}\in\mathbb{R}^{3\times H\times W}$.
Combining these components, we construct our proposed garment-invariant representation $\mathbf{I}_{gi}\in\mathbb{R}^{6\times H\times W}$ as follows:
(1) $\mathbf{I}_{gi}=\mathbf{I}_{vm}\oplus\mathbf{I}_{shm}$
where $\oplus$ denotes channel-wise concatenation.
Figure 5. Garment-invariant representation extraction from a person image.
3.3. Reliable Human Body Semantic Map Estimation
Unlike (Wu et al., 2025), which employs a simplified DensePose map as the human body semantic representation, our approach utilizes an even more simplified version, in which continuous UV coordinates are removed, as illustrated in Figure 6. This quantization enhances the learnability of the mapping from a garment-invariant representation to the human body semantic map. Following (Wu et al., 2025), we merged the lower and upper torso regions in the DensePose map. This was motivated by the observation that the boundary between these regions is not invariant to garment type, even for tight-fitting garments, as demonstrated in Figure 11.
((a))
((b))
((c)) (Wu et al., 2025)
((d)) Ours
Figure 6. Visual illustration of human body semantic maps: (a) The input image; (b) The DensePose map; (c) The human body semantic map of (Wu et al., 2025); (d) The human body semantic map of ours.
We obtain the human body semantic map through a two-step process: First, we extract the garment-invariant representation from the input image. Next, we transform this garment-invariant representation into the human body semantic map using a per-person BodyMap network.
To generate the training data for the BodyMap network tailored to a specific person, we record a video of the person wearing a tight-fitting garment while performing predefined movements, as described in (Wu et al., 2025). This person will later serve as the human model for per-garment dataset collection.
As illustrated in Figure 3, we use the recorded video to generate paired garment-invariant representations and human body semantic maps to train a BodyMap network, whose architecture follows pix2pixHD (Wang et al., 2018). The BodyMap network is then used to estimate the human body semantic map of the same person wearing a loose-fitting garment.
The BodyMap network is trained on a per-person basis and does not generalize across individuals. While this necessitates collecting a dedicated dataset for each human model, the process remains accessible, as it does not rely on specialized sensors to capture the body’s semantic map beneath loose-fitting garments. Moreover, since we typically use the same human model to collect multiple per-garment datasets, the trained BodyMap network can be efficiently reused.
3.4. Dataset Generation for Loose-Fitting Garments
For each garment item, we collect a garment-specific dataset, hereafter referred to as a per-garment dataset, to enable the training of garment-specific networks. Similar to the procedure in (Wu et al., 2025), we record a video of the human model wearing the loose-fitting garment performing predefined movements. In addition, we record a video of the same human model wearing an arbitrary tight-fitting garment and performing the same predefined movements, which is used to train the BodyMap network.
As illustrated in Figure 7, given a frame $\mathbf{I}$ of the loose-fitting garment videos, we obtain the garment image $\mathbf{I}_{g}$ and garment mask $\mathbf{I}_{m}$ using SAM2 (Ravi et al., 2024). Then, we obtain the virtual measurement garment image $\mathbf{I}_{vm}$, which is also used to construct the garment-invariant representation $\mathbf{I}_{gi}$. By inputting the garment-invariant representation into the trained BodyMap network, we obtain the human body semantic map $\mathbf{I}_{sdp}$ robustly under loose-fitting garments.
Figure 7. Per-garment dataset generation from recorded video frames of the human model. Unlike (Wu et al., 2025) that directly obtain the human body semantic map using DensePose (Güler et al., 2018), we extract garment-invariant representation and then transform it into the semantic map by our trained BodyMap network (red dashed box).
After processing each frame of the loose-fitting garment video to obtain its corresponding $\mathbf{I}_{g}$, $\mathbf{I}_{m}$, $\mathbf{I}_{vm}$, and $\mathbf{I}_{sdp}$, we finish generating the per-garment dataset for the loose-fitting garment. We preserve the sequential order of each frame in the dataset to utilize temporal information for training.
3.5. Training Pipeline of Virtual Try-On
We follow (Wu et al., 2025) to utilize hybrid person representation as the intermediate representation for garment synthesis. The hybrid person representation $\mathbf{I}_{hybrid}\in\mathbb{R}^{6\times H\times W}$ is a combination of virtual measurement garment $\mathbf{I}_{vm}$ and human body semantic map $\mathbf{I}_{sdp}$, and can be formally represented as:
(2) $\mathbf{I}_{hybrid}=\mathbf{I}_{vm}\oplus\mathbf{I}_{sdp}$
where $\mathbf{I}_{vm}$ captures the 3D pose of the human body which is crucial for $360^{\circ}$ virtual try-on, $\mathbf{I}_{sdp}$ provides pixel-level alignment guidance in 2D image space.
Unlike (Wu et al., 2025), which trains the garment synthesis network on a per-frame basis, we train a Recurrent Garment Synthesis (ReGarSyn) network with sequential data to enhance the temporal consistency of the try-on results. We insert a lightweight ConvLSTM (Shi et al., 2015) module into pix2pixHD (Wang et al., 2018) to obtain our ReGarSyn network.
As illustrated in Figure 13, given an input sequence $\{\mathbf{I}_{hybrid}^{1},\allowbreak\mathbf{I}_{hybrid}^{2},\allowbreak\ldots,\allowbreak\mathbf{I}_{hybrid}^{N}\}$ and a corresponding ground truth output sequence $\{(\mathbf{I}_{g}^{1},\allowbreak\mathbf{I}_{m}^{1}),\allowbreak(\mathbf{I}_{g}^{2},\allowbreak\mathbf{I}_{m}^{2}),\allowbreak\ldots,\allowbreak(\mathbf{I}_{g}^{N},\mathbf{I}_{m}^{N})\}$, the input and output of the ReGarSyn network at time step $t$ can be represented as:
(3) $\tilde{\mathbf{I}}_{g}^{t},\tilde{\mathbf{I}}_{m}^{t},\mathcal{C}^{t},\mathcal{H}^{t}=ReGarSyn(\mathbf{I}_{hybrid}^{t},\mathcal{C}^{t-1},\mathcal{H}^{t-1})$
Here, $\tilde{\mathbf{I}}_{g}^{t}$ and $\tilde{\mathbf{I}}_{m}^{t}$ denote the synthesized garment image and mask at time step $t$, respectively. $\mathcal{C}$ and $\mathcal{H}$ represent the cell state and hidden state of the ConvLSTM module, capturing temporal dependencies across the sequence.
The loss function $\mathcal{L}$ can be defined as:
(4) $\mathcal{L}=\sum_{t=1}^{N}PerFrame(\tilde{\mathbf{I}}_{g}^{t}\oplus\tilde{\mathbf{I}}_{m}^{t},\mathbf{I}_{g}^{t}\oplus\mathbf{I}_{m}^{t})$
where $PerFrame(\cdot,\cdot)$ denotes the per-frame loss function from pix2pixHD (Wang et al., 2018). Explicit temporal consistency loss functions are not employed, as the supervision provided by temporally consistent ground truth data inherently ensures temporal coherence.
3.6. Inference Pipeline of Virtual Try-On
During inference, we assume the user is wearing a tight-fitting garment, which allows us to efficiently extract a semantic map of the human body using DensePose (Güler et al., 2018). This assumption is practical and user-friendly, as it avoids the need for specialized wearable devices required by a prior approach (Chong et al., 2021).
As illustrated in Figure 10, given an input frame $\mathbf{I}^{t}$ at time step $t$, we extract the virtual measurement garment $\mathbf{I}_{vm}^{t}$ using BEV (Sun et al., 2022), and obtain the human body semantic map $\mathbf{I}_{sdp}^{t}$ via DensePose (Güler et al., 2018). Both processes are computationally efficient and suitable for real-time applications. Subsequently, as described in Equation 3, we input them together with $\mathcal{C}^{t-1}$ and $\mathcal{H}^{t-1}$ from the previous time step to the trained ReGarSyn network to obtain the synthesized garment image $\tilde{\mathbf{I}}_{g}^{t}$ and mask $\tilde{\mathbf{I}}_{m}^{t}$. Finally, we composite them onto the input frame to obtain the try-on result $\mathbf{I}_{o}^{t}$:
(5) $\mathbf{I}_{o}^{t}=\mathbf{I}^{t}\odot(\mathbf{1}-\tilde{\mathbf{I}}_{m}^{t})+\tilde{\mathbf{I}}_{g}^{t}\odot\tilde{\mathbf{I}}_{m}^{t}$
where $\odot$ denotes element-wise multiplication.
Our recurrent framework facilitates temporally coherent virtual try-on outputs over arbitrarily long input sequences while maintaining real-time performance and a constant memory footprint.
4. Experiments
4.1. Experiment Setting
Per-garment datastes. We collected per-garment datasets for four loose-fitting garments, including a hanfu, a down jacket, a hanbok, and a dress, using the same human model. Each dataset contains approximately 3,000 images. Additionally, we recorded a video of the same human model wearing a tight-fitting garment and performing similar movements for evaluation purposes.
Implementation details. We set the resolution of the BodyMap network to $512\times 384$ and the ReGarSyn networks to $576\times 432$. For each garment, we trained a per-garment model using the corresponding per-garment dataset. We trained each model on an NVIDIA RTX 4090 GPU for $40$ epochs, with a learning rate of $2\times 10^{-4}$. During training, we randomly sampled clips ranging in length from 8 to 60 frames. We use the Adam optimizer with $\beta_{1}=0.5$ and $\beta_{2}=0.999$. The training duration for a garment-specific model was approximately $12$ hours.
Comparison setting. We conduct both qualitative and quantitative evaluations of our method, using the prior per-garment virtual try-on approach (Wu et al., 2025) as the baseline for comparison. Additionally, we compare our results with OOTDiffusion (Xu et al., 2024b) and ViViD (Fang et al., 2024), which represent state-of-the-art image-based and video-based virtual try-on methods, respectively. Due to mismatches in input and output formats, retraining OOTDiffusion and ViViD on our datasets is not feasible, limiting the fairness of direct comparisons. Nonetheless, we include these methods to demonstrate that their reliance on a universal network for generalization across all garment types proves inadequate in practice, especially for rarely seen garments.
4.2. Qualitative Evaluation
We present a qualitative comparison of image quality among OOTDiffusion (Xu et al., 2024b), ViViD (Fang et al., 2024), the baseline method (Wu et al., 2025), and our method in Figure 8. Both OOTDiffusion and ViViD produce poor results for these loose-fitting garments, which supports our assertion that training a universal network on frequently seen garments cannot generalize well to all types of garments.
In the results produced by the baseline method, noticeable flaws appear in regions with loose-fitting garments—such as the wide sleeves of the hanfu (first row), the skirt hems of the hanbok (third row), and the hems of the dress (fourth row). These artifacts support our assertion that the baseline method struggles with discrepancies in input data distribution between the training and inference stages due to occlusions introduced by loose-fitting garments. While the down jacket (second row) does not exhibit such artifacts due to its relatively tight-fitting structure, it reveals temporal inconsistencies attributable to the baseline’s frame-by-frame garment synthesis strategy, as further evidenced by the quantitative results in Section 4.3. For a more intuitive video comparison, please refer to the supplementary video.
Figure 8. Qualitative comparison of our method against OOTDiffusion (Xu et al., 2024b), ViViD (Fang et al., 2024), and the baseline method (Wu et al., 2025).
Furthermore, to demonstrate the generalization ability of our ReGarSyn networks trained on per-garment datasets collected by a single human model, we present the virtual try-on results for unseen body shapes in Figure 14. This figure indicates that our strategy of using only one human model for each per-garment dataset is sufficient, as the ReGarSyn network successfully generalizes to a wide range of body shapes.
4.3. Quantitative Evaluation
Since collecting paired ground truth data for loose-fitting garment virtual try-on is impractical, we adopt evaluation metrics that do not rely on such data. Specifically, we use Kernel Inception Distance (KID) and Fréchet Inception Distance (FID) to assess image quality. To evaluate both image quality and temporal consistency, we employ Video Fréchet Inception Distance (VFID), which leverages the I3D backbone (Carreira and Zisserman, 2017). These metrics are computed using two videos: one showcasing the generated virtual try-on result and the other depicting the same person wearing the actual loose-fitting garment. Although the videos exhibit similar human motion, they are not strictly synchronized.
We report the quantitative results for four loose-fitting garments individually in Table 1. As shown in the table, both OOTDiffusion (Xu et al., 2024b) and ViViD (Fang et al., 2024) perform significantly worse than the baseline and our proposed method, which is consistent with the qualitative observations discussed in Section 4.2. Although the baseline occasionally surpasses our method in certain image quality metrics, our approach achieves comparable scores overall and consistently outperforms all others in VFID across all garments. This highlights the strength of our method in maintaining both temporal consistency and visual fidelity.
Table 1. Quantitative comparison with other approaches and ablation study on the ConvLSTM module.
Method Hanfu Down jacket Hanbok Dress
[…]
OOTDiffusion (Xu et al., 2024b) 0.073 77.27 1.138 0.180 154.43 1.234 0.257 187.30 1.359 0.116 100.47 1.036
ViViD (Fang et al., 2024) 0.190 153.88 1.266 0.200 173.29 1.115 0.439 285.40 1.254 0.125 105.57 0.996
Baseline (Wu et al., 2025) 0.043 51.25 1.041 0.102 97.81 1.027 0.022 28.85 0.942 0.071 66.75 0.992
Ours w/o ConvLSTM 0.037 44.24 1.027 0.102 98.41 1.029 0.020 26.92 0.933 0.072 67.00 0.981
Ours 0.034 42.88 1.025 0.099 96.86 1.024 0.021 27.01 0.930 0.073 67.26 0.980
4.4. Frame Rate Comparison
We compare the frame rate of our method against OOTDiffusion (Xu et al., 2024b), ViViD (Fang et al., 2024), the baseline approach (Wu et al., 2025), and a variant of our model that shares the same architecture but excludes the ConvLSTM module and is trained on a per-frame basis (”ours w/o ConvLSTM”). This variant is included to evaluate the impact of the ConvLSTM module on frame rate performance. All frame rates are measured using an NVIDIA RTX 4090 GPU.
As shown in Table 2, both OOTDiffusion and ViViD exhibit frame rates that fall short of real-time requirements. The baseline achieves the highest frame rate, while the ”ours w/o ConvLSTM” variant, which incorporates additional processing for the DensePose map, performs slightly slower. Our proposed method integrates a ConvLSTM module, resulting in a $14\%$ reduction in frame rate. Despite this, it maintains a frame rate of 10.5 frames per second, which remains adequate for real-time use.
Table 2. Comparison of frame rates. The introduction of the ConvLSTM module results in a slight decrease in frame rate but remains sufficient for real-time performance use.
Method Frame rate (fps) $\uparrow$
OOTDiffusion (Xu et al., 2024b) 0.32
ViViD (Fang et al., 2024) 1.32
Baseline (Wu et al., 2025) 12.38
Ours w/o ConvLSTM 12.15
Ours 10.50
4.5. Ablation Study on Garment-Invariant Representation
To demonstrate the necessity of our proposed garment-invariant representation and its components, we conduct ablation studies with the following alternative approaches for human body semantic map estimation:
  • DP: Directly using DensePose (Güler et al., 2018) to obtain the human body semantic map without any intermediate representation. This is to demonstrate the benefit of our proposed garment-invariant representation.
  • HM: Using heatmaps without removing any joints as the intermediate representation (Figure 9(a)). This is to show the significance of removing garment-sensitive joints.
  • SHM: Using heatmaps without garment-sensitive joints as the intermediate representation (Figure 9(b)). This reveals the importance of the virtual measurement garment (Wu et al., 2024).
  • VM: Using virtual measurement garment (Wu et al., 2024) as the intermediate representation (Figure 9(c)). This demonstrates the importance of joint heatmaps.
((a)) HM
((b)) SHM
((c)) VM
Figure 9. Three alternative intermediate representations used in ablation study for human body semantic map estimation.
We lack ground truth data for estimating human body semantic maps under loose-fitting garments, as obtaining paired images of the same person in identical poses wearing both tight-fitting and loose-fitting clothing is not feasible. Therefore, we can only conduct a qualitative comparison. The metrics used in Section 4.3 are unsuitable here since they cannot accurately reflect how well the semantic maps align with the human body.
As shown in Figure 15, our proposed garment-invariant representation enables reliable human body semantic map estimation. The semantic maps estimated by DensePose suffer from occlusion, e.g., the lower body region is missing due to the occlusion of the hanbok skirt (first row, DP column). Using joint heatmaps that include garment-sensitive joints results in inflated semantic maps (HM column). Utilizing joint heatmaps without garment-sensitive joints fails to provide sufficient information for semantic map estimation, resulting in distorted and incomplete results (SHM column). Relying solely on the virtual measurement garments does not provide pixel-level guidance for semantic map estimation, resulting in limbs being placed in incorrect regions (second row, VM column).
4.6. Ablation Study on ConvLSTM Module
We evaluate the effectiveness of the ConvLSTM module in enhancing temporal consistency through both qualitative and quantitative comparisons with a variant without ConvLSTM (”ours w/o ConvLSTM”), as described in Section 4.4.
To better illustrate the presence of jittering artifacts, we visualize the absolute differences between consecutive output frames using heatmaps, as shown in Figure 16. These visualizations reveal that the ”ours w/o ConvLSTM” variant exhibits noticeable jittering artifacts, particularly along the boundary of the loose skirt, which tends to vary in appearance even when the human remains relatively static. In contrast, our method, which incorporates the ConvLSTM module, achieves notably improved temporal consistency across frames. For a more intuitive comparison, please refer to the supplementary video.
5. Limitations and Future Work
In addition to the common limitations mentioned by (Wu et al., 2025), such as the necessity of data capture for each garment, the original garment is not removed, and inaccurate composition masks, we would like to highlight two limitations and their future work as follows:
Inability to convey actual size and fit.:
While our method generates visually plausible virtual try-on results for a wide range of body shapes and provides useful styling references (Figure 14), it does not validate the suitability of garment sizing due to the lack of actual user body measurements. This limitation is common among most existing image-based virtual try-on approaches. A promising future direction is to develop a garment synthesis framework that incorporates user-provided body dimensions to enable more accurate and personalized virtual try-on experiences.
Limitation in modeling transient physical dynamics.:
Although our recurrent framework incorporates temporal information by conditioning garment synthesis on preceding frames, its primary objective is to enforce inter-frame consistency rather than to model the transient physical dynamics of loose-fitting garments. The datasets we collected using the method of (Wu et al., 2025) comprise sequences of humans rotating in place, offering inadequate motion diversity for capturing the complex physical behaviors of loose garments. To address this limitation, a novel dataset collection method that introduces a broader range of motion while remaining accessible needs to be explored.
6. Conclusion
We introduce a novel per-garment approach that is the first to enable real-time and temporally coherent virtual try-on for loose-fitting garments. To enhance the quality of synthesized images, we propose a garment-invariant representation that facilitates accurate estimation of human body semantic maps under loose-fitting garments during per-garment dataset generation. This representation effectively mitigates the distribution gap between the training and inference stages of our garment synthesis network, resulting in significantly improved visual fidelity. To ensure temporal consistency, we propose a recurrent framework that leverages temporal dependencies for garment synthesis. Our inference pipeline demonstrates sufficient efficiency to support real-time virtual try-on and can process input frames of arbitrary length. Through extensive qualitative and quantitative evaluations, we demonstrate that our method outperforms existing approaches in both image quality and temporal consistency. We also conduct ablation studies to validate the effectiveness of the garment-invariant representation and the recurrent synthesis framework.